OpenStudy (anonymous):

Medal For Help :0) Which expression gives the solution of -5+2x2=6x? a. 2+ 4-(4)(6)(-5) =12 b.-5 + 25-(4)(2)(6) =-10 c. -6+ 36-(4)(2)(-5) =4 d. 6+ 36-(4)(2)(5) =4

OpenStudy (anonymous):

@shiraz14

OpenStudy (anonymous):

@destany9093

OpenStudy (shiraz14):

Are you sure you typed these options correctly? Because I think the last option should read -5 (instead of 5) and there's a square root sign missing in all the options?

OpenStudy (anonymous):

What are the approximate solutions of 2x2 -x +10=0? a.-2,2.5 b.-1.97, 2.47 c. -2.5,2 d.np solution

OpenStudy (anonymous):

@shiraz14 yes the 1 st example is correct

OpenStudy (shiraz14):

The answer to your second question is d. Actually the solution is imaginary

OpenStudy (anonymous):

yeap I got d too.

OpenStudy (shiraz14):

As for your first solution, I do believe that you meant to include square root signs in all the functions? Please clarify if this was included in the original question.

OpenStudy (anonymous):

Which method is the best method for solving the equation 8x2-13x+3=0? A.Square Roots B.Factoring C.Graphing D.Quadratic Formula

OpenStudy (shiraz14):

D. quadratic formula.

OpenStudy (anonymous):

#1 yes Square couldn't find the symbols

OpenStudy (anonymous):

How many solutions are there for 5x2+7x-4=0? A. 0 B.1 C. 2 D.3

OpenStudy (shiraz14):

For your level, the answer would be ... can you use the quadratic formula to work out the discriminant?

OpenStudy (anonymous):

Last question is c.2

OpenStudy (anonymous):

x=1 10 (−7±129 x=0.43578166916,−1.83578166916

OpenStudy (shiraz14):

You're correct

OpenStudy (anonymous):

TY I do know how to figure out alot but these few types and some graphs I'm working hard to learn

OpenStudy (anonymous):

THe perimeter of a rectangle is 54 cm. The area of the same rectangle is 176 cm2.What are the dimensions of the rectangle ?

OpenStudy (anonymous):

11 cm by 16 cm 8cm by 22 cm 5.5 cm by 32 cm 4cm by 44 cm

OpenStudy (shiraz14):

Let the length be x and the width be y. The perimeter is thus 2x + 2y = 54, i.e. x + y = 27cm. The area is xy = 176cm^2. Can you then try solving these set of simultaneous equations?

OpenStudy (anonymous):

kk going try

OpenStudy (shiraz14):

Alright, I'd need to log off now. Goodbye.

OpenStudy (anonymous):

ty again

OpenStudy (shiraz14):

You're welcome